OpenStudy (anonymous):

what is the integral of 0 to pi of cos(x+pi)dx

OpenStudy (anonymous):

is equal to\[-\int\limits_{0}^{\pi}\cos x dx\]

OpenStudy (anonymous):

0

OpenStudy (ash2326):

\[\int_{0}^{\pi} cos (x+\pi) dx\] integration of cos x= sin x so we get \[sin(x+\pi)\] \[sin(2\pi)-sin(\pi)\] so \[0-0\] 0